# Topology of Discrete Quantum Feedback Control

### v2

@@ IV.2 Eigenspectrum of a quantum channel under the periodic and open boundary conditions
In Fig. 6, we show the eigenspectrum of a quantum channel for the chiral Maxwell’s demon under the periodic boundary condition (PBC). Beside the highly degenerate zero eigenvalues $\xi_{n\neq 0}(k)$, the loop structure formed by the eigenvalue $\xi_{0}(k)$ is evident. The winding number is $w(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. The eigenvalue $\xi_{0}(k)$ is determined from the probability distribution of the position after feedback control as shown in Eq. (43). In fact, the eigenvalue $\xi_{0}(k)$ is nothing but the characteristic function of the probability distribution $\{p(r)\}$. The chiral transport caused by the feedback control increases the probability of a particle being found at site $r>0$ and makes the probability distribution $\{p(r)\}$ asymmetric with respect to $r>0$ and $r<0$. Thus, the winding number of the quantum channel is closely related to the chiral transport achieved by the feedback in analogy with the chiral edge state in the quantum Hall effect and the Hatano-Nelson model for non-Hermitian topological phases [33, 84, 85, 86, 87].
@@
Figure 6: Eigenspectrum of a quantum channel for the chiral Maxwell’s demon. A blue curve and a blue dot constitute the eigenspectrum under the periodic boundary condition for infinite system size, and orange dots show the eigenspectrum under the open boundary condition for system size $L=50$. Energy is measured in units of $J=1$, and we take $V=+\infty$ for simplicity. The duration of a feedback operation is $\tau=1$.
@@
Nontrivial topology around a point gap under the PBC implies the emergence of a non-Hermitian skin effect under the open boundary condition (OBC) [33, 72, 73, 88, 38, 74]. Orange dots in Fig. 6 show the eigenspectrum of the quantum channel under the OBC, which is drastically different from that under the PBC. Such an extraordinary sensitivity of the eigenspectrum to the boundary condition is a salient feature of the non-Hermitian skin effect. The steady state of the quantum channel under the PBC is uniform in space, while it is localized near the right edge of the system under the OBC [see Fig. 7(a) and (b)]. In Fig. 7(c) and (d), we show typical behavior of right and left eigenmodes under the OBC. The localization of the right and left eigenmodes at the opposite boundaries of the system indeed confirms the non-Hermitian skin effect, which is a consequence of the chiral transport induced by the chiral Maxwell’s demon.
@@
As seen from Fig. 7(c) and (d), the localization length of a right eigenmode is, in general, different from that of the corresponding left eigenmode due to the difference between the dynamics near the right boundary and those near the left boundary. Specifically, the right eigenmode in Fig. 7(c) is less localized than the left eigenmode in Fig. 7(d) because of the reflection at the right boundary. This phenomenon arises from the fact that the quantum channel under the OBC is defined from the time evolution under the same boundary condition so that the complete positivity and trace preservation are ensured (see also Sec. VIII.2).
@@
We note that the localization of right and left eigenmodes $\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}$ at opposite edges implies that the overlap $\langle\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}\rangle$ decreases exponentially as a function of the system size. From the eigemode expansion (14), this inditates that the expansion coeffecient $c_{n}$ becomes exponentially large, which significantly affects the relaxation dynamics of the system towards the steady state under repeated applications of a quantum channel [89, 90]. Specifically, if the non-Hermitian skin effect takes place, the relaxation time under the OBC generically diverges in the infinite-size limit [90].
@@
Figure 7: Magnitude $|\rho_{i,j}|$ of matrix elements of an eigenmode $\hat{\rho}_{n}=\sum_{i,j}\rho_{i,j}\ket{i}\bra{j}$ | plotted in the position basis for the system size of $L=20$ with the hopping amplitude $J=1$, the height of the potential $V=+\infty$, and the duration $\tau=1$ of feedback control. Steady states of the chiral Maxwell’s demon under the PBC and the OBC are shown in (a) and (b), respectively. Panels (c) and (d) show typical right and left eigenmodes under the OBC with eigenvalue 0.334 (see Fig. 6).
@@ IV.3 Topological transition
The chiral Maxwell’s demon described in Sec. IV.1 exhibits topological transitions as the duration $\tau$ of feedback control is changed. In Fig. 8, we show the eigenspectrum under the PBC for four different values of the duration of feedback control. As the duration of feedback control is increased, the absolute value of the winding number also increases. In Fig. 9(a), we show the dependence of the winding number on the duration of feedback control. The topological transitions occur at an almost constant rate as a function of the duration of feedback control.
Figure 8: Eigenspectrum of the quantum channel for the chiral Maxwell’s demon for four different values of the duration of the feedback. A blue curve and a blue dot constitute the eigenspectrum under the PBC for infinite system size, and orange dots show the eigenspectrum under the OBC for the system size of $L=50$. Energy is measured in units of $J=1$, and the height of the potential is set to $V=+\infty$. The duration of feedback control is (a) $\tau=0.5$, (b) $\tau=2$, (c) $\tau=3$, and (d) $\tau=3.9$. The winding number $w$ around the origin is (a) $w=0$, (b) $w=-2$, (c) $w=-3$, and (d) $w=-4$.
Figure 9: (a) Dependence of the winding number $w(\xi_{\mathrm{PG}}=0)$ on the duration $\tau$ of feedback control. (b) Velocity of a particle during feedback control [see Eq. (47)]. The height of the potential is set to be $V=+\infty$.
The topological transitions can physically be understood from the ballistic transport of a particle during the feedback control. In Fig. 9(b), we show the mean velocity
@@
of a particle during the feedback control. As seen from the figure, the particle shows ballistic transport at a constant velocity after an initial transient dynamics. To understand the relation between the ballistic tranport and the topological transitions, we rewrite the winding number (46) by using $z=e^{-ik}$ as
@@
is the probability-generating function of $\{p(r)\}$, and the integration countor in Eq. (48) is taken in the counterclockwise direction. From Eq. (48) and by using the argument principle, we have
@@
where $N_{\mathrm{p}}$ ($N_{0}$) is the number of poles (zeros) of the function $\xi_{0}(z)$ in $|z|<1$. The expression (50) gives a physical insight into topological feedback control. Let $\tilde{p}(r):=p(r-1)$ be a shifted probability distribution. Then, its probability-generating function is given by $z\xi_{0}(z)$. It follows then from Eq. (50) that the winding number is changed by $-1$ because the number of zeros is increased by one. The constant increase of the winding number can be understood from the ballistic transport of a particle, since the winding number is changed by $-1$ if the probability distribution of the displacement after feedback control is shifted by one site.
@@ IV.4 Robustness of chiral transport against disorder
The unidirectional transport induced by the chiral Maxwell’s demon is of a topological origin and therefore robust against external perturbations. Here we demonstrate that the topological feedback control enhances the stability of particle transport. By way of illustration, we consider a task of transporting a paricle from the leftmost site to the rightmost site in an $L$-site chain by using the chiral Maxwell’s demon. Suppose that the feedback Hamiltonian $\hat{H}_{j}$ is perturbed by onsite disorder potentials as
@@
where $W_{i}\in\mathbb{R}$ denotes a random potential. For comparison, we consider a non-topological feedback control with a quantum channel
@@
$\hat{H}_{1}^{\prime}=$ $0,$ (54b)
i.e., only the rightmost site is measured projectively and the dynamics is stopped when a particle is found at that site.
@@
with $\hat{x}=\sum_{j}j\ket{j}\bra{j}$ | and an initial state $\hat{\rho}_{\mathrm{i}}=\ket{1}\bra{1}$ | as a function of the number $N$ of feedback cycles. As shown in the figure, the chiral transport due to topological feedback control is robust against disoder as long as $|W_{i}|\ll V$, whereas the particle transport in the non-topological case is quickly suppressed due to the Anderson localization [91] in the presence of disorder. Thus, topological feedback control can enhance the stability of particle transport against disorder.
Figure 10: Expectation value of the position as a function of the number of feedback cycles. Red (black) dots show the result for topological (non-topological) feedback control. The parameters are set to be $L=20$, $J=1$, $\tau=1$, and $V=10$. The disorder potential $W_{i}$ at each site is randomly sampled from a uniform distribution in $[-W,W]$. We take $W=0$ in panel (a) and $W=1$ in panel (b). The results are averaged over 100 samples of disorder realization.
@@ V Symmetry classification of feedback control
Symmetry enriches topological phases since two topological phases may be distiguished in the presence of symmetry [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we discuss symmetry of quantum channels for discrete quantum feedback control. In Sec. V.1, we discuss unitary symmetry. In Sec. V.2, we consider antiunitary symmetry associated with every quantum channel. While the contents in these two subsections can be found in literature [92, 93, 33, 94, 66, 67], we reproduce them to make the present paper self-contained. In Sec. V.3, we discuss the Bernard-LeClair (BL) symmetry classes of quantum channels for discrete quantum feedback control. In particular, we find that only a subset of the BL symmetry classes are compatible with projective measurement. Consequently, we obtain the ten-fold symmetry classification of discrete quantum feedback control, as listed in Table 2.
@@ V.1 Unitary symmetry
We first consider unitary symmetry of quantum channels. Similarly to the case of open quantum systems described by the Lindblad equation [92, 93], a unitary symmetry of a quantum channel can be either weak symmetry or strong symmetry [94]. A CPTP map $\mathcal{E}$ has unitary symmetry if it commutes with a unitary superoperator $\mathcal{V}$, that is,
@@
where $\tilde{\mathcal{V}}$ is the matrix representation of $\mathcal{V}$ and given by a unitary operator acting on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. We call unitary symmetry in Eqs. (56) and (57) as weak symmetry [92, 93, 94]. If a CPTP map has weak symmetry, it can be block-diagonalized according to the symmetry eigenvalues. For example, the translation symmetry in Eq. (25) is a weak symmetry, and the argument $\vec{k}$ of the Bloch matrix $X(\vec{k})$ corresponds to the symmetry eigenvalue (see Appendix A.2).
Suppose that every Kraus operator of a CPTP map commutes with a unitary operator $\hat{V}$ as $[\hat{K}_{m},\hat{V}]=0\ (\forall m)$. Then, the CPTP map satisfies Eq. (57) with $\tilde{\mathcal{V}}=\hat{V}\otimes\hat{I}$ or $\tilde{\mathcal{V}}=\hat{I}\otimes\hat{V}^{*}$. Such symmetry is called strong symmetry [92, 93, 94], which gives a more stringent condition than weak symmetry. Strong symmetry is related to a conserved quantity of the dynamics. To see this, we consider the case of continuous symmetry and assume that every Kraus operator $\hat{K}_{m}$ commutes with $\hat{V}_{\theta}=\exp[-i\theta\hat{A}]$ for all $\theta\in\mathbb{R}$, where $\hat{A}$ is a Hermitian operator. This assumption implies $[\hat{K}_{m},\hat{A}]=0\ (\forall m)$. Then, for an arbitrary density matrix $\hat{\rho}$, we have
@@ V.2 Modular conjugation symmetry
A CPTP map has an antiunitary symmetry due to its Hermiticity-preserving nature [33]. To see this, we define the modular conjugation [95, 66] by an antiunitary operator $\tilde{J}$ acting on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$ as
@@
where we use the definition ~|$\tilde{\mathcal{E}}\ket{\hat{\rho}}=\ket{\mathcal{E}(\hat{\rho})}$ of the matrix representation [see Eq. (197)], the property of the modular conjugation [Eq. (60)], and the Hermiticity-preserving condition of a CPTP map [Eq. (16)]. Since Eq. (61) holds for arbitrary |$\ket{\hat{\rho}}$, we obtain
@@
V.3 Ten-fold symmetry classification of feedback control with projective measurement
The topological classification of non-Hermitian matrices is performed on the basis of the Bernard-LeClair (BL) symmetry classes [75], which are the non-Hermitian generalization [34, 35] of the Altland-Zirnbauer (AZ) symmetry classes [96]. Here, we discuss the BL symmetry classes of quantum channels for discrete quantum feedback control. In general, a symmetry of a CPTP map $\mathcal{E}$ that belongs to a BL symmetry class has one of the following representations [68]:

### v3

@@ IV.2 Eigenspectrum of a quantum channel under the periodic and open boundary conditions
In Fig. 6, we show the eigenspectrum of a quantum channel for the chiral Maxwell demon under the PBC. Besides the highly degenerate zero eigenvalues $\xi_{n\neq 0}(k)$, the loop structure formed by the eigenvalue $\xi_{0}(k)$ is evident. The winding number is $w(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. The eigenvalue $\xi_{0}(k)$ is determined from the probability distribution of the position after feedback control as shown in Eq. (43). In fact, the eigenvalue $\xi_{0}(k)$ is nothing but the characteristic function of the probability distribution $\{p(r)\}$. The chiral transport caused by feedback control increases the probability of a particle being found at site $r>0$ and makes the probability distribution $\{p(r)\}$ asymmetric with respect to $r>0$ and $r<0$. Thus, the winding number of the quantum channel is closely related to the chiral transport achieved by the feedback in analogy to the chiral edge state in the quantum Hall effect and the Hatano-Nelson model for non-Hermitian topological phases [33, 84, 85, 86, 87].
@@
Figure 6: Eigenspectrum of a quantum channel for the chiral Maxwell demon. A blue curve and a blue dot constitute the eigenspectrum under the periodic boundary condition for infinite system size, and orange dots show the eigenspectrum under the open boundary condition for system size $L=50$. Energy is measured in units of $J=1$, and we take $V=+\infty$ for simplicity. The duration of a feedback operation is $\tau=1$.
@@
Nontrivial topology around a point gap under the PBC implies the emergence of a non-Hermitian skin effect under the open boundary condition (OBC) [33, 72, 73, 88, 38, 74]. Orange dots in Fig. 6 show the eigenspectrum of the quantum channel under the OBC, which is drastically different from that under the PBC. Such an extraordinary sensitivity of the eigenspectrum to the boundary condition is a salient feature of the non-Hermitian skin effect. The steady state of the quantum channel under the PBC is uniform in space, while it is localized near the right edge of the system under the OBC [see Figs. 7(a) and 7(b)]. In Figs. 7(c) and 7(d), we show the typical behavior of right and left eigenmodes under the OBC. The localization of the right and left eigenmodes at the opposite boundaries of the system indeed confirms the non-Hermitian skin effect, which is a consequence of the chiral transport induced by the chiral Maxwell demon.
@@
As seen from Fig. 7(c) and 7(d), the localization length of a right eigenmode is, in general, different from that of the corresponding left eigenmode due to the difference between the dynamics near the right boundary and those near the left boundary. Specifically, the right eigenmode in Fig. 7(c) is less localized than the left eigenmode in Fig. 7(d) because of the reflection at the right boundary. This phenomenon arises from the fact that the quantum channel under the OBC is defined by the time evolution under the same boundary condition so that the complete positivity and trace preservation are ensured (see also Sec. VIII.2).
@@
We note that the localization of left and right eigenmodes $\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}$ at opposite edges implies that the overlap $\langle\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}\rangle$ decreases exponentially as a function of the system size. From the eigenmode expansion (14), this indicates that the expansion coefficient $c_{n}$ becomes exponentially large, which significantly affects the relaxation dynamics of the system towards the steady state under repeated applications of a quantum channel [89, 90]. Specifically, if the non-Hermitian skin effect takes place, the relaxation time under the OBC generically diverges in the infinite-size limit [90].
@@
Figure 7: Magnitude $|\rho_{i,j}|$ of matrix elements of an eigenmode $\hat{\rho}_{n}=\sum_{i,j}\rho_{i,j}\ket{i}\bra{j}$ | plotted in the position basis for a system size of $L=20$ with the hopping amplitude $J=1$, the height of the potential $V=+\infty$, and the duration $\tau=1$ of feedback control. Steady states of the chiral Maxwell demon under the PBC and the OBC are shown in panels (a) and (b), respectively. Panels (c) and (d) show typical right and left eigenmodes under the OBC with eigenvalue 0.334 (see Fig. 6).
@@ IV.3 Topological transition
The chiral Maxwell demon described in Sec. IV.1 exhibits topological transitions as the duration $\tau$ of feedback control is changed. In Fig. 8, we show the eigenspectrum under the PBC for four different values of the duration of feedback control. As the duration of feedback control is increased, the absolute value of the winding number also increases. In Fig. 9(a), we show the dependence of the winding number on the duration of feedback control. The topological transitions occur at an almost constant rate as a function of the duration of feedback control.
Figure 8: Eigenspectrum of the quantum channel for the chiral Maxwell demon for four different values of the duration of the feedback. A blue curve and a blue dot constitute the eigenspectrum under the PBC for infinite system size, and orange dots show the eigenspectrum under the OBC for a system size of $L=50$. Energy is measured in units of $J=1$, and the height of the potential is set to $V=+\infty$. The duration of feedback control is (a) $\tau=0.5$, (b) $\tau=2$, (c) $\tau=3$, and (d) $\tau=3.9$. The winding number $w$ around the origin is (a) $w=0$, (b) $w=-2$, (c) $w=-3$, and (d) $w=-4$.
Figure 9: (a) Dependence of the winding number $w(\xi_{\mathrm{PG}}=0)$ on the duration $\tau$ of feedback control. (b) Velocity of a particle during feedback control [see Eq. (47)]. The height of the potential is set to $V=+\infty$.
The topological transitions can physically be understood from the ballistic transport of a particle during feedback control. In Fig. 9(b), we show the mean velocity
@@
of a particle during feedback control. As seen from the figure, the particle shows ballistic transport at a constant velocity after an initial transient dynamics. To understand the relation between the ballistic transport and the topological transitions, we rewrite the winding number (46) by using $z=e^{-ik}$ as
@@
is the probability-generating function of $\{p(r)\}$, and the integration contour in Eq. (48) is taken in the counterclockwise direction. From Eq. (48) and by using the argument principle, we have
@@
where $N_{\mathrm{p}}$ ($N_{0}$) is the number of poles (zeros) of the function $\xi_{0}(z)$ in $|z|<1$. The expression (50) gives a physical insight into topological feedback control. Let $\tilde{p}(r):=p(r-1)$ be a shifted probability distribution. Then, its probability-generating function is given by $z\xi_{0}(z)$. It follows from Eq. (50) that the winding number is changed by $-1$ because the number of zeros is increased by $1$. The constant increase of the winding number can be understood from the ballistic transport of a particle since the winding number is changed by $-1$ if the probability distribution of the displacement after feedback control is shifted by one site.
@@ IV.4 Robustness of chiral transport against disorder
The unidirectional transport induced by the chiral Maxwell demon is of a topological origin and therefore robust against external perturbations. Here, we demonstrate that the topological feedback control enhances the stability of particle transport. By way of illustration, we consider a task of transporting a particle from the leftmost site to the rightmost site in an $L$-site chain by using the chiral Maxwell demon. Suppose that the feedback Hamiltonian $\hat{H}_{j}$ is perturbed by on-site disorder potentials as
@@
where $W_{i}\in\mathbb{R}$ denotes a random potential. For comparison, we consider a nontopological feedback control with a quantum channel
@@
$\hat{H}_{1}^{\prime}=$ $0;$ (54b)
i.e., only the rightmost site is measured projectively, and the dynamics stops when a particle is found at that site.
@@
with $\hat{x}=\sum_{j}j\ket{j}\bra{j}$ | and an initial state $\hat{\rho}_{\mathrm{i}}=\ket{1}\bra{1}$ | as a function of the number $N$ of feedback cycles. As shown in the figure, chiral transport due to topological feedback control is robust against disorder as long as $|W_{i}|\ll V$, whereas particle transport in the nontopological case is quickly suppressed due to Anderson localization [91] in the presence of disorder. Thus, topological feedback control can enhance the stability of particle transport against disorder.
Figure 10: Expectation value of the position as a function of the number of feedback cycles. Red (black) dots show the results for topological (nontopological) feedback control. The parameters are set as $L=20$, $J=1$, $\tau=1$, and $V=10$. The disorder potential $W_{i}$ at each site is randomly sampled from a uniform distribution in $[-W,W]$. We take $W=0$ in panel (a) and $W=1$ in panel (b). The results are averaged over 100 samples of disorder realization.
@@ V Symmetry classification of feedback control
Symmetry enriches topological phases since two topological phases may be distinguished in the presence of symmetry [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we discuss the symmetry of quantum channels for discrete quantum feedback control. In Sec. V.1, we discuss unitary symmetry. In Sec. V.2, we consider antiunitary symmetry associated with every quantum channel. While the content in these two subsections can be found in the literature [92, 93, 33, 94, 66, 67], we reproduce them to make the present paper self-contained. In Sec. V.3, we discuss the Bernard-LeClair (BL) symmetry classes of quantum channels for discrete quantum feedback control. In particular, we find that only a subset of the BL symmetry classes are compatible with projective measurements. Consequently, we obtain the tenfold symmetry classification of discrete quantum feedback control, as listed in Table 2.
@@ V.1 Unitary symmetry
We first consider unitary symmetry of quantum channels. Similarly to the case of open quantum systems described by the Lindblad equation [92, 93], unitary symmetry of a quantum channel can be either weak symmetry or strong symmetry [94]. A CPTP map $\mathcal{E}$ has unitary symmetry if it commutes with a unitary superoperator $\mathcal{V}$, that is,
@@
where $\tilde{\mathcal{V}}$ is the matrix representation of $\mathcal{V}$ which is given by a unitary operator acting on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. We call unitary symmetry in Eqs. (56) and (57) weak symmetry [92, 93, 94]. If a CPTP map has weak symmetry, it can be block diagonalized according to the symmetry eigenvalues. For example, translation symmetry in Eq. (25) is weak symmetry, and the argument $\vec{k}$ of the Bloch matrix $X(\vec{k})$ corresponds to the symmetry eigenvalue (see Appendix A.2).
Suppose that every Kraus operator of a CPTP map commutes with a unitary operator $\hat{V}$ as $[\hat{K}_{m},\hat{V}]=0\ (\forall m)$. Then, the CPTP map satisfies Eq. (57) with $\tilde{\mathcal{V}}=\hat{V}\otimes\hat{I}$ or $\tilde{\mathcal{V}}=\hat{I}\otimes\hat{V}^{*}$. Such symmetry is called strong symmetry [92, 93, 94], which gives a more stringent condition than weak symmetry. Strong symmetry is related to a conserved quantity of the dynamics. To see this relation, we consider the case of continuous symmetry and assume that every Kraus operator $\hat{K}_{m}$ commutes with $\hat{V}_{\theta}=\exp[-i\theta\hat{A}]$ for all $\theta\in\mathbb{R}$, where $\hat{A}$ is a Hermitian operator. This assumption implies $[\hat{K}_{m},\hat{A}]=0\ (\forall m)$. Then, for an arbitrary density matrix $\hat{\rho}$, we have
@@ V.2 Modular conjugation symmetry
A CPTP map has an antiunitary symmetry due to its Hermiticity-preserving nature [33]. To see this property, we define the modular conjugation [95, 66] by an antiunitary operator $\tilde{J}$ acting on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$ as
@@
where we use the definition ~|$\tilde{\mathcal{E}}\ket{\hat{\rho}}=\ket{\mathcal{E}(\hat{\rho})}$ of the matrix representation [see Eq. (197)], the property of the modular conjugation [Eq. (60)], and the Hermiticity-preserving property of a CPTP map [Eq. (16)]. Since Eq. (61) holds for arbitrary |$\ket{\hat{\rho}}$, we obtain
@@
V.3 Tenfold symmetry classification of feedback control with projective measurement
The topological classification of non-Hermitian matrices is performed on the basis of the BL symmetry classes [75], which are the non-Hermitian generalizations [34, 35] of the Altland-Zirnbauer (AZ) symmetry classes [96]. Here, we discuss the BL symmetry classes of quantum channels for discrete quantum feedback control. In general, the symmetry of a CPTP map $\mathcal{E}$ that belongs to a BL symmetry class has one of the following representations [68]: